A man has a collection of 29 classic cars and motorcycles. The total number of wheels for both the cars and motorcycles is 96. How many cars does this man have? The answer is 19 cars and 10 motorcycles. What are the steps taken to solve the equation.

Answer by JimboP1977(311)

C+M=29
4C+2M=96
(96-4C)/2=M
(48-2C)=M
C+48-2C=29
-C=-19
C=19
M=10